OpenStudy (anonymous):

Simplify. Assume that all expressions are defined. (1/x)+(1/2x)/((x+4)/(x-2))

OpenStudy (anonymous):

hiiiiiii

OpenStudy (nathan917):

I don't get it. sorry :/

OpenStudy (anonymous):

\[\frac{ \frac{ 1}{x }+\frac{ 1 }{ 2x} }{ \frac{ x+4 }{ x-2 }}\] how bout now @nathan917

OpenStudy (anonymous):

does it seem confusing still

OpenStudy (anonymous):

its ok if u cant do it

OpenStudy (anonymous):



OpenStudy (anonymous):

@Eyad

OpenStudy (anonymous):